# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v2

@@ 4 Methods
This section describes the Bayesian and Frequentist estimation frameworks used for parameter inference, uncertainty quantification, and forecasting. Let $Y=(y_{t_{1}},\ldots,y_{t_{n}})$ denote the observed data at discrete time points $t_{1},\ldots,t_{n}$, and let $\boldsymbol{\theta}$ represent the vector of unknown model parameters to be estimated. For the Lotka–Volterra (LV) model, $\boldsymbol{\theta}=(\alpha,\beta,\gamma,\delta)$. For the generalized logistic model (GLM), $\boldsymbol{\theta}=(r,p,K)$. For the SEIUR model, $\boldsymbol{\theta}=(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$. Both estimation approaches assume a normal error structure to ensure comparability of results.
@@ 4.1 Bayesian Inference
Bayesian inference integrates prior knowledge with observed data to obtain the posterior distribution of model parameters [87]. This probabilistic framework is particularly effective when prior information is available or when data are sparse, noisy, or subject to partial observability, as it provides comprehensive uncertainty quantification through posterior distributions [88, 89].
@@ 4.1.1 Bayes’ Rule
According to Bayes’ theorem, the posterior distribution of parameters given the data is proportional to the product of the prior distribution and the likelihood:
$p(\boldsymbol{\theta}\mid Y)\propto p(\boldsymbol{\theta})\,p(Y\mid\boldsymbol{\theta}),$ (7)
where $p(\boldsymbol{\theta})$ is the prior distribution encoding existing knowledge about the parameters before observing data, $p(Y\mid\boldsymbol{\theta})$ is the likelihood function representing the probability of observing the data given parameter values, and $p(\boldsymbol{\theta}\mid Y)$ is the posterior distribution of parameters after incorporating the observed data.
@@ 4.1.2 Likelihood Function
We assume that observation errors are independent and identically distributed (i.i.d.) following a normal distribution with constant variance. For the Lotka–Volterra model with observed prey and predator populations, the likelihood is:
$y_{t_{j}}^{\text{prey}}\mid\boldsymbol{\theta}\sim\mathcal{N}(x_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{prey}}^{2}),\quad y_{t_{j}}^{\text{predator}}\mid\boldsymbol{\theta}\sim\mathcal{N}(y_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{predator}}^{2}),$ (8)
independently for $j=1,\ldots,n$, where $x_{\boldsymbol{\theta}}(t)$ and $y_{\boldsymbol{\theta}}(t)$ denote the prey and predator population solutions to the Lotka–Volterra ODE system given parameters $\boldsymbol{\theta}$, and $\sigma_{\text{prey}}^{2}$ and $\sigma_{\text{predator}}^{2}$ are observation error variances.
@@
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\frac{dC_{\boldsymbol{\theta}}(t_{j})}{dt},\sigma^{2}\right),\quad j=1,\ldots,n,$ (9)
where $C_{\boldsymbol{\theta}}(t)$ is the solution to the GLM ODE given parameters $\boldsymbol{\theta}$.
@@
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\kappa\rho E_{\boldsymbol{\theta}}(t_{j}),\sigma^{2}\right),\quad j=1,\ldots,n,$ (10)
@@ 4.1.3 Prior Distributions
Prior distributions encode existing knowledge or beliefs about parameter values before observing data [44] . For parameters with established biological or epidemiological interpretations, we specify informative priors based on previous studies. When prior knowledge is limited, we use weakly informative or uniform priors over plausible parameter ranges. For structural identifiability analysis with simulated data, priors were centered on true parameter values with varying degrees of precision to assess the impact of prior strength on parameter recovery. Table 28 summarizes the prior distributions used for each model. The normal error variance $\sigma^{2}$ is assigned a weakly informative inverse-gamma or half-Cauchy prior to allow the data to inform the observation error scale.
@@ 4.1.4 Posterior Sampling and Convergence Diagnostics
Posterior distributions are approximated using Markov Chain Monte Carlo (MCMC) sampling implemented through the Hamiltonian Monte Carlo (HMC) algorithm in Stan [62, 65]. For each model, we ran four independent chains with a sufficient number of iterations (typically 2,000–4,000) after a burn-in period to ensure convergence to the stationary distribution. Convergence was assessed using the Gelman–Rubin diagnostic $\hat{R}$ statistic [65], which compares between-chain and within-chain variance; values of $\hat{R}<1.05$ (ideally $\hat{R}\approx 1.01$) indicate successful convergence. Effective sample size (ESS) was monitored to ensure adequate sampling of the posterior distribution. Trace plots and posterior density plots were visually inspected to confirm mixing and convergence.
Posterior summaries are reported as medians with 95% prediction intervals (PIs), which represent the central 95% of the posterior distribution. For forecasting, we propagate parameter uncertainty by sampling parameters values from the posterior distribution, solving the ODE for each sample, and summarizing the resulting forecast distribution with median and 95% prediction intervals (CIs).
@@ 4.1.5 Computational Implementation
Bayesian inference was conducted using the (BFF) toolbox [26], an R package designed for fitting and forecasting ODE-based epidemic models. The BFF toolbox provides an automated workflow that generates Stan code based on user-specified model structures, priors, and data inputs, eliminating the need for users to program directly in Stan. The toolbox outputs include posterior parameter distributions, convergence diagnostics ($\hat{R}$, ESS), trace plots, posterior density plots, and forecasting results with quantified uncertainty. Performance metrics including mean absolute error (MAE), mean squared error (MSE), weighted interval score (WIS), and 95% PI coverage are automatically computed to evaluate model fit and forecast accuracy.
@@ 4.2.1 Parameter Estimation
$\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}}{\text{argmin}}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\boldsymbol{\theta})\right)^{2},$ (11)
where $\mu_{j}(\boldsymbol{\theta})$ is the model-predicted value at time $t_{j}$ given parameters $\boldsymbol{\theta}$. For the Lotka–Volterra model, $\mu_{j}(\boldsymbol{\theta})$ consists of predicted prey and predator populations $(x_{\boldsymbol{\theta}}(t_{j}),y_{\boldsymbol{\theta}}(t_{j}))$. For the GLM, $\mu_{j}(\boldsymbol{\theta})=dC_{\boldsymbol{\theta}}(t_{j})/dt$. For the SEIUR model, $\mu_{j}(\boldsymbol{\theta})=\kappa\rho E_{\boldsymbol{\theta}}(t_{j})$.
In our study, the Frequentist inference is performed using MATLAB’s fmincon with the Sequential Quadratic Programming (SQP) algorithm. To mitigate sensitivity to initial conditions and reduce the risk of convergence to local minima, the optimization is wrapped within a MultiStart framework using a CustomStartPointSet, with an optional RandomStartPointSet.
The observed data variance is estimated as:
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\hat{\boldsymbol{\theta}})\right)^{2}.$ (12)
4.2.2 Uncertainty Quantification via Parametric Bootstrap
To quantify parameter uncertainty and construct prediction intervals, we employ a parametric bootstrap procedure [33, 35]. This approach generates synthetic datasets by resampling from the fitted model and re-estimates parameters for each synthetic dataset. The distribution of bootstrap parameter estimates characterizes sampling variability. The steps are as follows:
  1. 1.
@@ 4.2.2 Uncertainty Quantification via Parametric Bootstrap
    $y_{t_{j}}^{b}\sim\mathcal{N}\left(\mu_{j}(\hat{\boldsymbol{\theta}}),\hat{\sigma}^{2}\right),\quad j=1,\ldots,n,$ (13)
@@
  2. 2.
    Re-estimate parameters: For each bootstrap sample $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$, solve the NLS optimization problem (Eq. 11) to obtain bootstrap parameter estimates $\hat{\boldsymbol{\theta}}^{b}$.
  3. 3.
    Construct prediction intervals: Use the empirical distribution of $\{\hat{\boldsymbol{\theta}}^{b}:b=1,\ldots,B\}$ to construct prediction intervals. For example, the 95% prediction interval for parameter $\theta_{k}$ is given by the 2.5th and 97.5th percentiles of $\{\hat{\theta}_{k}^{b}:b=1,\ldots,B\}$.
For forecasting, we propagate parameter uncertainty by solving the ODE model for each bootstrap parameter estimate $\hat{\boldsymbol{\theta}}^{b}$ to obtain forecast trajectories $\hat{y}^{b}(t_{n+h})$ for $h$-step-ahead predictions. The 95% prediction interval for $y(t_{n+h})$ is constructed from the 2.5th and 97.5th percentiles of $\{\hat{y}^{b}(t_{n+h}):b=1,\ldots,B\}$.
@@ 4.2.3 Computational Implementation
Frequentist inference was implemented using the QuantDiffForecast (QDF) MATLAB toolbox [25], which provides a comprehensive framework for parameter estimation, uncertainty quantification, and forecasting for ODE models. The QDF toolbox supports multiple optimization algorithms, flexible error structures (normal, Poisson, negative binomial), user-defined ODE systems, and automated bootstrap-based uncertainty quantification. The toolbox outputs parameter estimates with prediction intervals, fitted model trajectories, forecast distributions, and performance metrics (MAE, MSE, WIS, 95% PI coverage).
@@ 4.3 Performance Metrics
4.3.1 Mean Absolute Error (MAE)
@@
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right|.$ (14)
MAE provides a direct measure of forecast accuracy on the original data scale. Lower MAE values indicate better point forecast performance. MAE is less sensitive to outliers compared to MSE.
4.3.2 Mean Squared Error (MSE)
MSE measures the average squared deviation between model predictions and observed data:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right)^{2}.$ (15)
MSE penalizes larger errors more heavily than MAE due to the squaring operation, making it more sensitive to outliers. Lower MSE values indicate better fit. The square root of MSE (RMSE) is sometimes reported to return the metric to the original data scale.
4.3.3 Coverage of the 95% prediction interval
The 95% PI coverage quantifies the proportion of observed data points that fall within the 95% prediction interval, providing a measure of uncertainty calibration:
$\text{95\% PI Coverage}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(L_{t_{i}}<y_{t_{i}}<U_{t_{i}}),$ (16)
where $L_{t_{i}}$ and $U_{t_{i}}$ are the lower and upper bounds of the 95% prediction interval at time $t_{i}$, and $\mathbf{1}(\cdot)$ is an indicator function equal to 1 if the condition is true and 0 otherwise. Ideally, the coverage should be close to 95%. Coverage significantly below 95% indicates overly narrow intervals (underestimation of uncertainty), while coverage above 95% suggests overly wide intervals (overestimation of uncertainty).
4.3.4 Weighted Interval Score (WIS)
WIS is a proper scoring rule that evaluates the quality of the entire predictive distribution by combining sharpness (interval width) and calibration (penalties for observations outside intervals) [74, 73]. The interval score (IS) for a single prediction interval at level $\alpha$ is:
$\text{IS}_{\alpha}(F,y)=(u-l)+\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)+\frac{2}{\alpha}(y-u)\mathbf{1}(y>u),$ (17)
where $l$ and $u$ represent the $\frac{\alpha}{2}$ and $\left(1-\frac{\alpha}{2}\right)$ quantiles of the forecast distribution $F$, respectively. The IS consists of three components:
  • Sharpness: (ul)(u-l), the width of the central $(1-\alpha)\times 100\%$ prediction interval. Narrower intervals receive lower scores, rewarding precise forecasts.
@@ 4.3.4 Weighted Interval Score (WIS)
  • Overprediction penalty: $\frac{2}{\alpha}(y-u)\mathbf{1}(y>u)$, penalizes observations exceeding the upper bound $u$, with the penalty proportional to the distance (yu)(y-u).
To comprehensively evaluate the full predictive distribution, we compute the WIS by averaging interval scores over multiple interval levels $(1-\alpha_{1})<(1-\alpha_{2})<\cdots<(1-\alpha_{K})$ along with the predictive median $\tilde{y}$ (which can be viewed as an interval at level $(1-\alpha_{0})\to 0$):
$\text{WIS}_{\alpha_{0:K}}(F,y)=\frac{1}{K+\frac{1}{2}}\left(w_{0}|y-\tilde{y}|+\sum_{k=1}^{K}w_{k}\,\text{IS}_{\alpha_{k}}(F,y)\right),$ (18)
where $w_{k}=\frac{\alpha_{k}}{2}$ for $k=1,\ldots,K$ and $w_{0}=\frac{1}{2}$. WIS quantifies how close the entire predictive distribution is to the observed data in units on the original data scale. Lower WIS values indicate better forecast performance, balancing sharpness and calibration.

### v3

@@ 4 Methods
This section describes the Bayesian and frequentist estimation frameworks used for parameter inference, uncertainty quantification, and forecasting. Let $Y=(y_{t_{1}},\ldots,y_{t_{n}})$ denote the observed data at discrete time points $t_{1},\ldots,t_{n}$, and let $\boldsymbol{\theta}$ represent the vector of unknown model parameters to be estimated. For the LV model, $\boldsymbol{\theta}=(\alpha,\beta,\gamma,\delta)$. For the GLM, $\boldsymbol{\theta}=(r,p,K)$. For the SEIUR model, $\boldsymbol{\theta}=(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$. Both estimation approaches assume a normal error structure to ensure the comparability of results.
@@ 4.1 Bayesian inference
Bayesian inference integrates prior knowledge with observed data to obtain the posterior distribution of the model’s parameters [82]. This probabilistic framework is particularly effective when prior information is available or when data are sparse, noisy, or subject to partial observability, as it provides comprehensive uncertainty quantification through posterior distributions [83].
@@ 4.1.1 Bayes’ rule
According to Bayes’ theorem, the posterior distribution of the parameters, given the data, is proportional to the product of the prior distribution and the likelihood:
$p(\boldsymbol{\theta}\mid Y)\propto p(\boldsymbol{\theta})\,p(Y\mid\boldsymbol{\theta}),$ (4.1)
where $p(\boldsymbol{\theta})$ is the prior distribution encoding existing knowledge about the parameters before observing the data, $p(Y\mid\boldsymbol{\theta})$ is the likelihood function representing the probability of observing the data given the parameter values, and $p(\boldsymbol{\theta}\mid Y)$ is the posterior distribution of parameters after incorporating the observed data.
@@ 4.1.2 Likelihood function
We assume that observation errors are independent and identically distributed (i.i.d.) following a normal distribution with constant variance. For the LV model with observed prey and predator populations, the likelihood is:
$y_{t_{j}}^{\text{prey}}\mid\boldsymbol{\theta}\sim\mathcal{N}(x_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{prey}}^{2}),\quad y_{t_{j}}^{\text{predator}}\mid\boldsymbol{\theta}\sim\mathcal{N}(y_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{predator}}^{2}),$ (4.2)
independently for $j=1,\ldots,n$, where $x_{\boldsymbol{\theta}}(t)$ and $y_{\boldsymbol{\theta}}(t)$ denote the prey and predator population solutions to the LV ODE system given the parameters $\boldsymbol{\theta}$, and $\sigma_{\text{prey}}^{2}$ and $\sigma_{\text{predator}}^{2}$ are the observation error variances.
@@
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\frac{dC_{\boldsymbol{\theta}}(t_{j})}{dt},\sigma^{2}\right),\quad j=1,\ldots,n,$ (4.3)
where $C_{\boldsymbol{\theta}}(t)$ is the solution to the GLM ODE given the parameters $\boldsymbol{\theta}$.
@@
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\kappa\rho E_{\boldsymbol{\theta}}(t_{j}),\sigma^{2}\right),\quad j=1,\ldots,n,$ (4.4)
@@ 4.1.3 Prior distributions
Prior distributions encode the existing knowledge or beliefs about parameter values before observing the data [44]. For parameters with established biological or epidemiological interpretations, we specify informative priors based on previous studies. When prior knowledge is limited, we use weakly informative or uniform priors over plausible parameter ranges. For structural identifiability analyses with simulated data, the priors were centered on true parameter values with varying degrees of precision to assess the impact of the prior strength on parameter recovery. Tables 1323, and 18 summarizes the prior distributions used for LV model. The normal error variance $\sigma^{2}$ is assigned a weakly informative inverse-gamma or half-Cauchy prior to allow the data to inform the observation error scale.
@@ 4.1.4 Posterior sampling and convergence diagnostics
Posterior distributions are approximated using MCMC sampling implemented through the Hamiltonian Monte Carlo (HMC) algorithm in Stan [61, 64]. For each model, we ran four independent chains with a sufficient number of iterations (typically 2000–4000) after a burn-in period to ensure convergence to the stationary distribution. Convergence was assessed using the Gelman–Rubin diagnostic $\hat{R}$ statistic [64], which compares between-chain and within-chain variance; values of $\hat{R}<1.05$ (ideally $\hat{R}\approx 1.01$) indicate successful convergence. Effective sample size (ESS) was monitored to ensure adequate sampling of the posterior distribution. Trace plots and posterior density plots were visually inspected to confirm mixing and convergence.
Posterior summaries are reported as medians with 95% PIs, which represent the central 95% of the posterior distribution. For forecasting, we propagate parameter uncertainty by sampling parameters values from the posterior distribution, solving the ODE for each sample, and summarizing the resulting forecast distribution with the median and 95% PIs.
@@ 4.1.5 Computational implementation
Bayesian inference was conducted using the BFF toolbox [26], an R package designed for fitting and forecasting ODE-based epidemic models. The BFF toolbox provides an automated workflow that generates Stan code based on user-specified model structures, priors, and data inputs, eliminating the need for users to program directly in Stan. The toolbox outputs include posterior parameter distributions, convergence diagnostics ($\hat{R}$, ESS), trace plots, posterior density plots, and forecasting results with quantified uncertainty. Performance metrics including MAE, MSE, WIS, and 95% PI coverage are automatically computed to evaluate model fit and forecast accuracy.
@@ 4.2.1 Parameter estimation
$\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}}{\text{argmin}}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\boldsymbol{\theta})\right)^{2},$ (4.5)
where $\mu_{j}(\boldsymbol{\theta})$ is the model-predicted value at time $t_{j}$ given parameters $\boldsymbol{\theta}$. For the LV model, $\mu_{j}(\boldsymbol{\theta})$ consists of predicted prey and predator populations $(x_{\boldsymbol{\theta}}(t_{j}),y_{\boldsymbol{\theta}}(t_{j}))$. For the GLM, $\mu_{j}(\boldsymbol{\theta})=dC_{\boldsymbol{\theta}}(t_{j})/dt$. For the SEIUR model, $\mu_{j}(\boldsymbol{\theta})=\kappa\rho E_{\boldsymbol{\theta}}(t_{j})$.
In our study, the frequentist inference is performed using MATLAB’s fmincon with the sequential quadratic programming (SQP) algorithm. To mitigate sensitivity to the initial conditions and reduce the risk of convergence to local minima, the optimization is wrapped within a MultiStart framework using a CustomStartPointSet, with an optional RandomStartPointSet.
The observed data variance is estimated as follows:
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\hat{\boldsymbol{\theta}})\right)^{2}.$ (4.6)
4.2.2 Uncertainty quantification via parametric bootstraping
To quantify parameter uncertainty and construct PIs, we use a parametric bootstrap procedure [33, 35]. This approach generates synthetic datasets by resampling from the fitted model and re-estimates the parameters for each synthetic dataset. The distribution of bootstrap parameter estimates characterizes the sampling variability. The steps are as follows.
  1. 1)
@@ 4.2.2 Uncertainty quantification via parametric bootstraping
    $y_{t_{j}}^{b}\sim\mathcal{N}\left(\mu_{j}(\hat{\boldsymbol{\theta}}),\hat{\sigma}^{2}\right),\quad j=1,\ldots,n,$ (4.7)
@@
  2. 2)
    Re-estimate parameters: For each bootstrap sample $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$, solve the NLS optimization problem (Eq 4.5) to obtain the bootstrap parameter estimates $\hat{\boldsymbol{\theta}}^{b}$.
  3. 3)
    Construct PIs: Use the empirical distribution of $\{\hat{\boldsymbol{\theta}}^{b}:b=1,\ldots,B\}$ to construct PIs. For example, the 95% PI for parameter $\theta_{k}$ is given by the 2.5th and 97.5th percentiles of $\{\hat{\theta}_{k}^{b}:b=1,\ldots,B\}$.
For forecasting, we propagate parameter uncertainty by solving the ODE model for each bootstrap parameter estimate $\hat{\boldsymbol{\theta}}^{b}$ to obtain the forecast trajectories $\hat{y}^{b}(t_{n+h})$ for $h$-step-ahead predictions. The 95% PI for $y(t_{n+h})$ is constructed from the 2.5th and 97.5th percentiles of $\{\hat{y}^{b}(t_{n+h}):b=1,\ldots,B\}$.
@@ 4.2.3 Computational implementation
Frequentist inference was implemented using the QuantDiffForecast (QDF) MATLAB toolbox [25], which provides a comprehensive framework for parameter estimation, uncertainty quantification, and forecasting for ODE models. The QDF toolbox supports multiple optimization algorithms, flexible error structures (normal, Poisson, negative binomial), user-defined ODE systems, and automated bootstrap-based uncertainty quantification. The toolbox outputs parameter estimates with PIs, fitted model trajectories, forecast distributions, and performance metrics (MAE, MSE, WIS, 95% PI coverage).
@@ 4.3 Performance metrics
4.3.1 Mean absolute error
@@
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right|.$ (4.8)
MAE provides a direct measure of forecast accuracy on the original data scale. Lower MAE values indicate better point forecast performance. MAE is less sensitive to outliers compared with MSE.
4.3.2 Mean squared error
MSE measures the average squared deviation between the model’s predictions and the observed data
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right)^{2}.$ (4.9)
MSE penalizes larger errors more heavily than MAE due to the squaring operation, making it more sensitive to outliers. Lower MSE values indicate a better fit. The square root of MSE (RMSE) is sometimes reported to return the metric to the original data scale.
4.3.3 Coverage of the 95% PI
The 95% PI coverage quantifies the proportion of observed data points that fall within the 95% PI, providing a measure of uncertainty calibration:
$\text{95\% PI Coverage}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(L_{t_{i}}<y_{t_{i}}<U_{t_{i}}),$ (4.10)
where $L_{t_{i}}$ and $U_{t_{i}}$ are the lower and upper bounds of the 95% PI at time $t_{i}$, and $\mathbf{1}(\cdot)$ is an indicator function equal to 1 if the condition is true and 0 otherwise. Ideally, the coverage should be close to 95%. Coverage significantly below 95% indicates overly narrow intervals (underestimation of uncertainty), while coverage above 95% suggests overly wide intervals (overestimation of uncertainty).
4.3.4 Weighted interval score
WIS is a proper scoring rule that evaluates the quality of the entire predictive distribution by combining sharpness (interval width) and calibration (penalties for observations outside intervals) [73, 72]. The interval score (IS) for a single PI at level $\alpha$ is:
$\text{IS}_{\alpha}(F,y)=(u-l)+\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)+\frac{2}{\alpha}(y-u)\mathbf{1}(y>u),$ (4.11)
where $l$ and $u$ represent the $\frac{\alpha}{2}$ and $\left(1-\frac{\alpha}{2}\right)$ quantiles of the forecast distribution $F$, respectively. The IS consists of the following three components:
  • Sharpness: (ul)(u-l), which is the width of the central $(1-\alpha)\times 100\%$ PI. Narrower intervals receive lower scores, rewarding precise forecasts.
@@ 4.3.4 Weighted interval score
  • Overprediction penalty: $\frac{2}{\alpha}(y-u)\mathbf{1}(y>u)$, which penalizes observations exceeding the upper bound $u$, with the penalty proportional to the distance (yu)(y-u).
To comprehensively evaluate the full predictive distribution, we compute the WIS by averaging the interval scores over multiple interval levels $(1-\alpha_{1})<(1-\alpha_{2})<\cdots<(1-\alpha_{K})$ along with the predictive median $\tilde{y}$ (which can be viewed as an interval at the level $(1-\alpha_{0})\to 0$):
$\text{WIS}_{\alpha_{0:K}}(F,y)=\frac{1}{K+\frac{1}{2}}\left(w_{0}|y-\tilde{y}|+\sum_{k=1}^{K}w_{k}\,\text{IS}_{\alpha_{k}}(F,y)\right),$ (4.12)
where $w_{k}=\frac{\alpha_{k}}{2}$ for $k=1,\ldots,K$ and $w_{0}=\frac{1}{2}$. The WIS quantifies how close the entire predictive distribution is to the observed data in units on the original data scale. Lower WIS values indicate better forecast performance, balancing sharpness and calibration.